a lot containing 8 tires components is sampled by a quality inspector; the lot contains 2 bad components. the inspector takes a sample of 4. find the expected value of the number of good components in the sample
Answer by ewatrrr(24785)

p(def) = 2/8 = 1/4, p(good) = 3/4
Sample of 4
E(good) = (3/4)4 = 3